Two six-sided dice are tossed. What is the probability that their sum is greater than 7, given that the first roll was a 4?
Answer by dabanfield(803)

There are 6*6 = 36 possibilities for outcomes of throwing the dice. 15 of these result in a sum greater than 7. So the probability of throwing greater than a 7 is 15/36= 5/12.
The first throw is irrelevant since it is an independent event.